# Simultaneously optimized mechanical stiffness and cell growth on scaffold by morphology learning

### v1

@@ Abstract
The morphology of microstructures in scaffolds is a critical factor influencing both mechanical properties and biological performance, such as cell growth, in bone tissue engineering. While metamaterials with optimized mechanical properties, such as hyperelasticity, thermal efficiency, or energy absorption, have been designed through complex microstructural geometry with the help of numerical simulation of multi-physics, optimizing biological performance alongside these physical traits remains a significant challenge. This is primarily because biological performance must be evaluated through laboratory experiments. To address this issue, we propose a novel approach that simultaneously optimizes mechanical stiffness and cell growth effectiveness through a data-driven morphology learning method. Our approach effectively extracts and learns shape patterns from a specifically selected dataset of microstructures recognized for their superior cell growth performance. We then integrate these patterns into a computational framework based on topology optimization, enabling the generation of new microstructures that maintain high levels of cell growth while enhancing mechanical stiffness. Using this morphology learning-based method, we identified a scaffold design for bone tissue engineering that demonstrates a significant improvement in both mechanical stiffness (by 29.69%) and cell proliferation (by 37.05% on Day 7 and 33.30% on Day 14).
@@
Scaffold, Structural pattern, Neural network, Cell culture
@@ Introduction
Figure 1: Achieving simultaneously optimized mechanical and biological performance. a Neural SDF-based morphology learning captures and embeds structural patterns from different datasets into network coefficients, enabling the conversion of latent codes into shapes with varying patterns represented by implicit SDF. b Seven distinct types of Triply Periodic Minimal Surfaces (TPMS) are used as the shape basis in our microstructure dataset for morphology learning, where each TPMS type is represented by an implicit basis function – allowing for the generation of intermediate shapes through interpolation. c A dataset filtered by a tissue engineering expert is used to generate a latent space of shape patterns that demonstrate good biological performance (the routine highlighted in blue arrows), where the latent space is then applied in a latent code-based topology optimization (TO) process to create structures with enhanced mechanical stiffness. As an alternative (displayed in the green routine), the latent space generated from an unfiltered (complete) dataset tends to result in structures with comparatively poorer biological performance. Random exploration within the latent space of biologically favorable shape patterns, without the guidance of finite element analysis (FEA) driven TO, may lead to structures with significantly weaker mechanical stiffness (as shown in the yellow routine).
@@
With the advancement of Artificial Intelligence (AI) techniques, a variety of AI-based topology optimization methods have been developed to either reduce or completely replace the need for iterative operations [35]. These methods can be broadly categorized into two types: 1) supervised approaches [36, 37], which primarily rely on image processing pipelines but often suffer from limited generalization capabilities, and 2) unsupervised approaches [38, 39, 40], where AI solvers optimize objectives by minimizing loss functions using the auto-differentiation capabilities of neural network-based computing. However, because gradient-based optimization necessitates explicitly defining optimization objectives as loss functions, directly incorporating biological performance into the optimization process remains a significant challenge.
@@ Results
As shown in the bottom row of Fig. 4a, this latent code-based interpolation effectively preserves the distinct shape patterns of $H_{a}$ and $H_{b}$ across the interpolation range, demonstrating the robustness of capturing shape patterns.
In addition to comparing the shapes, we also quantitatively analyzed the changes in surface area, volume, and the area-to-volume ratio between SDF interpolation and latent code-based interpolation. The results are plotted as curves in Fig. 4b. It is evident that direct SDF interpolation results in significant fluctuations in all these metrics, primarily due to the drastic changes in topology during the middle range of interpolation. In contrast, latent code-based interpolation produces much smoother transitions, reflecting a more consistent preservation of shape characteristics.
@@
In the first mechanical experiment, we evaluate the mechanical stiffness of the structures optimized by using the latent spaces (i.e., the networks) learned from both the full dataset and the filtered dataset. A structure is first randomly selected from the filtered dataset, which also exists in the full dataset (see the left-most one in Fig. 6a&b). The latent codes of this structures are also obtained from the two datasets to serve as the initial values for the computation of structure optimization. The objective function combines structural compliance and volume fraction, with five volume fractions used in the optimization as: 0.3, 0.35, 0.4, 0.45, and 0.5. The optimized structures are shown in Fig. 6a&b. The compliance values of the optimized structures are presented in Fig. 6c, where lower values indicate greater structural stiffness. Interestingly, the structures optimized within the shape space learned from the filtered dataset exhibit greater stiffness than those optimized within the space from the full dataset. This is due to the full dataset containing many structures with weak stiffness under the loading conditions defined in Fig. 5. As a result, the average performance of structures from the full dataset is slightly reduced compared to that of the filtered dataset. This experiment also highlights that careful selection of structures to include into the dataset is crucial. The convergence curves of compliance and volume fraction for the structure optimized using VF = 0.4 in the filtered dataset are shown in Fig. 6d and  6e. The bumps in these curves are caused by dynamic changes in the projection parameter. The volume fraction of the optimized structure converges well to 0.4, and compliance is effectively minimized.
@@
Figure 8: Mechanical compression tests of structures optimized by using the latent-code learned from the filter dataset vs. the conventional scaffold. a An initial structure with a volume fraction of 0.28. b Three compliance optimized structures are obtained with different target volume fractions (VF). The right-most shows the conventional scaffold with its volume fraction being about 0.45. The scaffold specimens with dimensions $12\mathrm{mm}\times 12\mathrm{mm}\times 18\mathrm{mm}$ that are 3D printed according to different structures shown in the first row. c The hardware setup employed in our mechanical compression tests to verify the mechanical stiffness of structures generated by our method via measuring the compression modulus. d Experimental results as stress-strain curves. e Bar-charts of the compression modulus values of each sample. f Bar-chart of the compression modulus values with standard derivation tested on 5 samples for each structure.
@@
Figure 9: Biological performance verification for the structures generated by the latent-code learned from the filter dataset. a The structures generated by optimizing samples randomly selected from the full dataset into a volume fraction as 0.45. b The structures generated by optimizing samples randomly selected from the filtered dataset – all resultant structures have VF=0.45. c The conventional scaffold design taken for comparison purpose. d The Alamar Blue assay was selected to assess and compare the cell proliferation across various structures, evaluating their biological performance. e The performance in hADSC proliferation results. f The averaged performance across the five structures for each dataset.
@@
Figure 10: Bio-images and cell proliferation of different scaffold designs. a Confocal microscopy images and b SEM images of hADSC on our optimized scaffold with VF=0.45 and the conventional scaffold at different time points, where red arrows represent the proliferated cells. Cell nuclei is stained blue (DAPI) and F-actin is stained green (Alexa Fluor 488) in confocal images. c hADSC and d Saos-2 cell viability and proliferation results on scaffolds with different volume fractures.
@@ Discussion
While the results from our experimental tests are promising, there is still room for improving the performance of our approach. In the current implementation, the dataset of structures with ‘good’ biological performance was constructed by excluding those identified as having ‘poor’ performance by a tissue engineering expert. This introduces potential bias stemming from the expert’s existing knowledge. A more rigorous selection process would involve laboratory testing of all structures in the dataset, retaining only those that demonstrate superior biological performance. This will potentially lead to structures with better biological performance. Another limitation is that the sample structures are exclusively derived from TPMS although the shape variation has been enriched by interpolating between seven different TPMS types. We anticipate that incorporating a wider range of structure types (e.g., those inspired by nature [57, 58]) could further increase the diversity of structural shapes and enhance biological performance.
From a broader perspective, this work introduces a general approach to integrating factors that cannot be directly evaluated through numerical simulations into structural topology optimization, which has focused primarily on maximizing mechanical stiffness. By leveraging the data-driven morphology learning methodology as proposed in this paper, we can simultaneously optimize multiple objectives even when they are difficult to formulate mathematically. These objectives may include not only mechanical performance but also factors such as aesthetics, physiological comfort, psychological well-being, and biological compatibility. This approach opens up many new possibilities for designing structures that balance diverse and complex requirements, making it applicable to a broader range of fields from biomedical engineering to product design. For example, psychological and aesthetic aspects play a significant role in consumer product design, which can now be potentially addressed alongside structural optimization.
@@ Methods
Our approach of structural optimization, which preserves the shape patterns extracted from morphology learning, builds upon the density-based topology optimization method SIMP. We first provide a brief introduction to SIMP. Key components of the SIMP framework will later be adapted for our pattern-preserving topology optimization.
@@
Confocal microscopy imaging was conducted to evaluate cell adhesion and spreading status on the scaffolds. On day 7 and 14 of the cell culture, cell-seeded scaffolds were fixed with 10% formalin (Sigma-Aldrich, UK), permeabilized with 0.1% Triton X-100 (Sigma-Aldrich, UK) in PBS, and stained with Alexa Fluor™ 488 Phalloidin (Invitrogen, USA) and 4’,6-Diamidino-2-phenylindole dihydrochloride (DAPI; Invitrogen, USA) according to manufacturer’s guidence. Confocal images were captured using a Leica SP8 LIGHTNING confocal microscope (Leica, Germany) with 405 nm and 499 nm lasers. Scanning electron microscopy (SEM) was used to further evaluate cell attachment and proliferation at days 7 and 14. The fixed cell-seeded scaffolds were dehydrated by gradient ethanol (50%, 70%, 80%, 90%, and 100%), followed by the treatment of 1:1 ethanol and hexamethyldisilazane (HMDS, Sigma-Aldrich, UK) and pure HMDS, considering 15 mins each step. The dehydrated samples were air-dried for volatilization, surface coated with 6 nm gold-palladium (4:1), and imaged using a TESCAN MIRA3 SEM (TESCAN, Czech) at an accelerating voltage of 2 kV.

### v3

# Simultaneous Topology Optimization of Differentiable and Non-Differentiable Objectives via Morphology Learning: Stiffness and Cell Growth on Scaffold

@@ Abstract
Topology optimization of microstructures plays a critical role in optimizing functional performance across diverse engineering applications. While metamaterials with enhanced mechanical properties – such as hyperelasticity, energy absorption, and thermal efficiency – are commonly designed using complex microstructural geometries and multi-physics simulations, achieving the simultaneous optimization of mechanical performance and non-differentiable objectives remains a significant challenge. In this work, we propose a novel framework for simultaneous topology optimization of differentiable and non-differentiable objectives via a data-driven morphology learning approach. The framework extracts shape patterns from a curated dataset of microstructures recognized for their superior performance in specific functional applications. To showcase the versatility of the approach, we apply it to the optimization of scaffolds for bone tissue engineering, with cell growth as a representative functional objective. By integrating learned morphology patterns into a topology optimization process, the method generates microstructures that effectively balance mechanical stiffness and biological performance, such as enhanced cell proliferation. As a case study, we demonstrate a scaffold design that improves mechanical stiffness by 29.69% and cell growth by 37.05% on Day 7 and 33.30% on Day 14. This approach highlights the general applicability of the proposed framework for optimizing a broad range of engineering challenges, beyond the specific case of cell growth.
@@
Structural pattern, Topology optimization, Neural network, Scaffold, Cell culture
@@ Introduction
Figure 1: Achieving simultaneously optimized mechanical and biological performance. a Neural SDF-based morphology learning captures and embeds structural patterns from different datasets into network coefficients, enabling the conversion of latent codes into shapes with varying patterns represented by implicit SDF. b Seven distinct types of Triply Periodic Minimal Surfaces (TPMS) are used as the shape basis in our microstructure dataset for morphology learning, where each TPMS type is represented by an implicit basis function – allowing for the generation of intermediate shapes through interpolation. c A dataset filtered by a tissue engineering expert is used to generate a latent space of shape patterns that demonstrate good biological performance (the routine highlighted in blue arrows), where the latent space is then applied in a latent code-based TO process to create structures with enhanced mechanical stiffness. As an alternative (displayed in the green routine), the latent space generated from an unfiltered (complete) dataset tends to result in structures with comparatively poorer biological performance. Random exploration within the latent space of biologically favorable shape patterns, without the guidance of finite element analysis (FEA) driven TO, may lead to structures with significantly weaker mechanical stiffness (as shown in the yellow routine).
@@
With the advancement of Artificial Intelligence (AI) techniques, a variety of AI-based TO methods have been developed to either reduce or completely replace the need for iterative operations [35]. These methods can be broadly categorized into two types: 1) supervised approaches [36, 37], which primarily rely on image processing pipelines but often suffer from limited generalization capabilities, and 2) unsupervised approaches [38, 39, 40], where AI solvers optimize objectives by minimizing loss functions using the auto-differentiation capabilities of neural network-based computing. However, because gradient-based optimization necessitates explicitly defining optimization objectives as loss functions, directly incorporating biological performance into the optimization process remains a significant challenge. With the advancement of AI techniques, a variety of AI-based TO methods have been developed to either reduce or completely replace the need for iterative operations [35].
@@ Results
As shown in the bottom row of Fig. 4a, this latent code-based interpolation effectively preserves the distinct shape patterns of $H_{a}$ and $H_{b}$ across the interpolation range, demonstrating the robustness of capturing shape patterns. In addition to comparing the shapes, we also quantitatively analyzed the changes in surface area, volume, and the area-to-volume ratio between SDF interpolation and latent code-based interpolation. The results are plotted as curves in Fig. 4b. It is evident that direct SDF interpolation results in significant fluctuations in all these metrics, primarily due to the drastic changes in topology during the middle range of interpolation. In contrast, latent code-based interpolation produces much smoother transitions, reflecting a more consistent preservation of shape characteristics.
We conducted an additional study to demonstrate the generalization capability of our latent code-based representation obtained through morphology learning. Specifically, as illustrated in Fig. 5, we utilized a neural network $f_{\theta}(\mathbf{x},\mathbf{z})$ trained on a TPMS dataset to approximate various unseen models. For a new model represented by the SDF function $s^{w}(\mathbf{x})$, its approximation is achieved by solving
${\mathbf{z}_{w}=\arg\min_{\mathbf{z}}\sum_{\mathbf{x}}|f_{\theta}(\mathbf{x},\mathbf{z})-s^{w}(\mathbf{x})|.}$ (3)
The resultant latent code $\mathbf{z}_{w}$ is determined through backpropagation, as illustrated in Fig. 3a, while keeping the network coefficients unchanged. The geometric approximation error is evaluated as the difference between $s^{w}(\mathbf{x})$ and $f_{\theta}(\mathbf{x},\mathbf{z}_{w})$ at every surface point $\mathbf{x}$ where $s^{w}(\mathbf{x})=0$. Notably, the approximation errors for the unseen TPMS model are significantly smaller compared to the other two unseen models. This is because the shape patterns learned and stored in $f_{\theta}(\cdot)$ are derived from the TPMS dataset, which exhibits substantial variations compared to the lattice and plate models.
Figure 5: Shape approximation by the latent code-based representation. The neural network $f_{\theta}(\mathbf{x},\mathbf{z})$ learned from a TPMS dataset is employed to approximate: an unseen TPMS model, a lattice structure, and a model of intersected plates in the top row. The geometric approximation errors are visualized by color maps on the surfaces of the reconstructed models shown in the bottom row.
@@
In the first mechanical experiment, we evaluate the mechanical stiffness of the structures optimized by using the latent spaces (i.e., the networks) learned from both the full dataset and the filtered dataset. A structure is first randomly selected from the filtered dataset, which also exists in the full dataset (see the left-most one in Fig. 7a&b). The latent codes of this structures are also obtained from the two datasets to serve as the initial values for the computation of structure optimization. The objective function combines structural compliance and volume fraction, with five volume fractions used in the optimization as: 0.3, 0.35, 0.4, 0.45, and 0.5. The optimized structures are shown in Fig. 7a&b. The compliance values of the optimized structures are presented in Fig. 7c, where lower values indicate greater structural stiffness. Interestingly, the structures optimized within the shape space learned from the filtered dataset exhibit greater stiffness than those optimized within the space from the full dataset. This is due to the full dataset containing many structures with weak stiffness under the loading conditions defined in Fig. 6. As a result, the average performance of structures from the full dataset is slightly reduced compared to that of the filtered dataset. This experiment also highlights that careful selection of structures to include into the dataset is crucial. The convergence curves of compliance and volume fraction for the structure optimized using VF = 0.4 in the filtered dataset are shown in Fig. 7d and  7e. The bumps in these curves are caused by dynamic changes in the projection parameter. The volume fraction of the optimized structure converges well to 0.4, and compliance is effectively minimized.
@@
Figure 9: Mechanical compression tests of structures optimized by using the latent-code learned from the filter dataset vs. the conventional scaffold. a An initial structure with a volume fraction of 0.28. b Three compliance optimized structures are obtained with different target volume fractions (VF). The right-most shows the conventional scaffold with its volume fraction being about 0.45. The scaffold specimens with dimensions $12\mathrm{mm}\times 12\mathrm{mm}\times 18\mathrm{mm}$ that are 3D printed according to different structures shown in the first row. c The hardware setup employed in our mechanical compression tests to verify the mechanical stiffness of structures generated by our method via measuring the compression modulus. d Experimental results as stress-strain curves. e Bar-charts of the compression modulus values of each sample. f Bar-chart of the compression modulus values with standard derivation tested on 5 samples for each structure. The significance levels are set as * p < 0.05, ** p < 0.01, and *** p < 0.001.
@@
Figure 10: Biological performance verification for the structures generated by the latent-code learned from the filter dataset. a The structures generated by optimizing samples randomly selected from the full dataset into a volume fraction as 0.45. b The structures generated by optimizing samples randomly selected from the filtered dataset – all resultant structures have VF=0.45. c The conventional scaffold design taken for comparison purpose. d The Alamar Blue assay was selected to assess and compare the cell proliferation across various structures, evaluating their biological performance. e The performance in hADSC proliferation results, where the measurements are taken on 5 specimens for each group – i.e., the bar-chart is generated by the mean values with standard derivation. f The averaged performance across the five structures for each dataset. The significance levels are set as * p < 0.05, ** p < 0.01, and *** p < 0.001.
Figure 11: Bio-images and cell proliferation of different scaffold designs. a Confocal microscopy images (×$\times 80$ magnification) and b SEM images (×$\times 500$ magnification) of hADSC on our optimized scaffold with VF=0.45 and the conventional scaffold at different time points, where red arrows represent the proliferated cells. Cell nuclei is stained blue (DAPI) and F-actin is stained green (Alexa Fluor 488) in confocal images. c hADSC and d Saos-2 cell viability and proliferation results on scaffolds with different volume fractures. The significance levels are set as * p < 0.05, ** p < 0.01, and *** p < 0.001.
@@ Discussion
While the results from our experimental tests are promising, there is still room for improving the performance of our approach. In the current implementation, the dataset of structures with ‘good’ biological performance was constructed by excluding those identified as having ‘poor’ performance by a tissue engineering expert. This introduces potential bias stemming from the expert’s existing knowledge. This may leave some structures with ‘poor’ biological performance in the selected dataset, potentially reducing the effectiveness of our approach. A more rigorous selection process would involve laboratory testing of all structures in the dataset to only retain those with superior biological performance in the selected dataset. This will potentially lead to structures with better biological performance. Another limitation is that the sample structures are exclusively derived from TPMS although the shape variation has been enriched by interpolating between seven different TPMS types. We anticipate that incorporating a wider range of structure types (e.g., those inspired by nature [57, 58]) could further increase the diversity of structural shapes and enhance biological performance.
The effectiveness of our approach has been verified on a unit microstructure with specified dimensions. This microstructure can be periodically replicated and merged to occupy a larger region, which can then be trimmed into a specific bone shape. These operations can be efficiently performed using Boolean operations on the resultant structures generated by our approach, as they are represented as implicit solids defined by $f_{\theta}(\mathbf{x})$. The feasibility of additively manufacturing such periodically applied structures over large regions has been demonstrated in prior research[43]. Additionally, when fabricating the scaffold using FDM, water-soluble materials can be employed for support structures when necessary.
From a broader perspective, this work introduces a general approach to integrating factors that cannot be directly evaluated through numerical simulations into structural TO, which has focused primarily on maximizing mechanical stiffness. By leveraging the data-driven morphology learning methodology as proposed in this paper, we can simultaneously optimize multiple objectives even when they are difficult to formulate mathematically. These objectives may include not only mechanical performance but also factors such as aesthetics, physiological comfort, psychological well-being, and biological compatibility. This approach opens up many new possibilities for designing structures that balance diverse and complex requirements, making it applicable to a broader range of fields from biomedical engineering to product design. For example, psychological and aesthetic aspects play a significant role in consumer product design, which can now be potentially addressed alongside structural optimization.
@@ Methods
Our approach of structural optimization, which preserves the shape patterns extracted from morphology learning, builds upon the density-based TO method SIMP. We first provide a brief introduction to SIMP. Key components of the SIMP framework will later be adapted for our pattern-preserving TO.
@@
Confocal microscopy imaging was conducted to evaluate cell adhesion and spreading status on the scaffolds. On day 7 and 14 of the cell culture, cell-seeded scaffolds were fixed with 10% formalin (Sigma-Aldrich, UK), permeabilized with 0.1% Triton X-100 (Sigma-Aldrich, UK) in PBS, and stained with Alexa Fluor™ 488 Phalloidin (Ex/Em = 490/525 nm, Invitrogen, USA) and 4’,6-Diamidino-2-phenylindole dihydrochloride (Ex/Em = 350/470 nm, DAPI; Invitrogen, USA) according to manufacturer’s guidence. Confocal images were captured using a Leica SP8 LIGHTNING confocal microscope (Leica, Germany) with 405 nm and 499 nm lasers. Scanning electron microscopy (SEM) was used to further evaluate cell attachment and proliferation at days 7 and 14. The fixed cell-seeded scaffolds were dehydrated by gradient ethanol (50%, 70%, 80%, 90%, and 100%), followed by the treatment of 1:1 ethanol and hexamethyldisilazane (HMDS, Sigma-Aldrich, UK) and pure HMDS, considering 15 mins each step. The dehydrated samples were air-dried for volatilization, surface coated with 6 nm gold-palladium (4:1), and imaged using a TESCAN MIRA3 SEM (TESCAN, Czech) at an accelerating voltage of 2 kV.